OpenStudy (anonymous):

Given ABC with vertices A(0, 8), B(3, 0) and C(-3, 0). Find the length of the altitude from A. (This is the second part of the question. The first part was to show that ABC was isosceles, which I did. But I am not sure how to solve this.) FYI (AC=AB=the root of 73 and BC=6) But how do I find this question.

OpenStudy (anonymous):

The box, next to ABC, means triangle. Sorry about the confusion...

OpenStudy (wolf1728):

From the graph it is easy to see that the height of altitude AD is 8. But if you want to solve this mathematically then Altitude AD² = AB² - DB² Altitude AD² = 73 - 9 Altitude AD = sqrt(64) Altitude AD = 8

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

But where does D come from?

OpenStudy (wolf1728):

I just entered the letter 'D' as the midpoint of line CB. I thought it showed the Pythagorean Theorem calculations more clearly.